# Hyperbolic Fracton Model, Subsystem Symmetry and Holography III: Extension to Generic Tessellations

### v1

@@ Abstract
We generalize the Hyperbolic Fracton Model from the $\{5,4\}$ tessellation to generic tessellations, and investigate its core properties: subsystem symmetries, fracton mobility, and holographic correspondence. While the model on the original tessellation has features reminiscent of the flat-space lattice cases, the generalized tessellations exhibit a far richer and more intricate structure. The ground-state degeneracy and subsystem symmetries are generated recursively layer-by-layer, through the inflation rule, but without a simple, uniform pattern. The fracton excitations follow exponential-in-distance and algebraic-in-lattice-size growing patterns when moving outward, and depend sensitively to the tessellation geometry, differing qualitatively from both type-I or type-II fracton model on flat lattices. Despite this increased complexity, the hallmark holographic features — subregion duality via Rindler reconstruction, the Ryu-Takayanagi formula for mutual information, and effective black hole entropy scaling with horizon area — remain valid. These results demonstrate that the holographic correspondence in fracton models persists in generic tessellations, and provide a natural platform to explore more intricate subsystem symmetries and fracton physics.
@@ I Introduction
The anti-de Sitter/conformal field theory (AdS/CFT) correspondence [1, 2, 3] is a breakthrough in theoretical physics, as it postulates a duality between a gravitational theory in a $(d+1)$-dimensional AdS spacetime and a $d$-dimensional conformal field theory (CFT) on its boundary. This holographic principle has reshaped our understanding of quantum gravity while also providing a powerful framework for studying strongly correlated systems [4, 5, 6, 7, 8]. In the meantime, the mathematical complexity of the original AdS/CFT arising from string theory has motivated the search for simpler toy models that capture the essential properties of holography [9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
In parallel, the discovery of fracton phases of matter unveiled a new frontier in many-body physics, expanding the established paradigms of topological order [19, 20, 21, 22, 23, 24, 25]. These fracton phases are characterized by quasiparticle excitations with restricted mobility and symmetries which act on lower-dimensional subsystems, sensitive to the geometry of the lattice beside topology [26, 27, 28]. Moreover, certain fracton theories share structural similarities with linearized gravity [29, 30, 31], which motivated the development of fracton models on hyperbolic lattices as toy models of holography [16, 17, 18].
This direction leads to a fertile ground to explore in hyperbolic lattices as they provide a discrete representation of AdS space [32]. Recent advances in crafting hyperbolic lattices in physical platforms, such as circuit QED [33, 34], have opened opportunities for the experimental simulation of quantum phenomena in hyperbolic space. The Hyperbolic Fracton Model (HFM) first introduced in Ref. [16, 17] leverages this synergy by defining a classical fracton model on the {5,4} tessellation, creating an exactly solvable model of holography. Recent works have solidified this connection by demonstrating a duality between the HFM and a $p$-adic model of AdS/CFT [35, 36].
@@
Here we generalize the original HFM to all $\{p,q\}$ tessellations of the hyperbolic plane, extending beyond the original simple cases, and discover much more complicated behavior without analogue in the type-I and type-II fracton models in flat 3D lattices [37]. The construction of our model is based on a set of geometry-dependent, recursive inflation rules that dictate the subsystem symmetries of the system. Through a detailed analysis of the model’s ground state degeneracy and properties, we derive scaling laws governed by the tessellations’ parameters $p$ and $q$. Another central result of our generalized model is the Rindler reconstruction, a procedure through which bulk spin configurations within a minimal wedge are uniquely determined by measurements on a corresponding boundary subregion. This provides a discrete analogue to the entanglement wedge reconstruction in continuum AdS/CFT. We also demonstrate that the mutual information between complementary boundary subregions, which is the classical analogue of entanglement entropy, follows a discrete form of the Ryu-Takayanagi formula [38], therefore linking boundary entanglement to the length of the shared boundary. Furthermore, we establish that the removal of bulk spins simulates the formation of a black hole. This induces a change in entropy proportional to the boundary area of the defect region, and offers a discrete analogue of the Bekenstein-Hawking entropy formula [39]. Finally, we have made quantitative analysis of how fractons multiply and grow exponentially in distance and algebraically in system size when being pushed away to further layers. This feature is qualitatively different from neither type-I or type-II fracton models on flat cubic lattice, and applies to quantum fracton orders defined on lattices embedded in 2D hyperbolic space $\times$ 1D flat space too [40].
@@
This paper is structured as follows: Section II provides a brief review of the essential concepts of the AdS/CFT correspondence that underpin our model. Section III elaborates on the recursive construction of generic hyperbolic tessellation geometry. Section IV introduces the Plaquette Ising model as a foundation for the hyperbolic generalization and presents the Hyperbolic Fracton Model [16, 17]. The following sections present our results. In Section V, we derive expressions for the ground state degeneracy and residual entropy. Then, the holographic properties of the model are elaborated through the Rindler reconstruction in Section VI, mutual information calculations in Section VII, and an analysis of black hole entropy in Section VIII. Stepping away from the ground-state properties, section IX is devoted to the analysis of fracton excitations. Finally we conclude our paper with a discussion in Section X.
@@ II AdS/CFT correspondence
The first concrete realization of the holographic principle is the AdS/CFT correspondence [1], which is a conjectured duality between a $(d+1)$-dimensional gravitational theory in an Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) on its boundary. AdS space is a maximally symmetric Lorentzian manifold with negative curvature, for example, AdS3 can be visualized as a temporal stack of two-dimensional hyperbolic planes $\mathbb{H}^{2}$. Meanwhile, a CFT is a field theory that is invariant under conformal symmetries, particularly powerful for describing systems at their critical points. The AdS/CFT duality is a strong-weak duality, where the boundary CFT that is strongly coupled and analytically intractable, corresponds to a weakly coupled gravitational theory in the bulk.
@@ III Hyperbolic Tessellations
A regular tessellation is a tiling of a surface by identical, regular polygonal faces, where the same number of faces meet at each vertex. These structures are uniquely classified by their Schläfli symbol $\{p,q\}$, where $p$ is the number of edges of each polygonal face, and $q$ is the number of faces meeting at each vertex of the tiling. The geometry of the embedding space is determined by the relationship between $p$ and $q$. For a flat Euclidean plane, tessellations must satisfy $1/p+1/q=1/2$, which permits only three configurations: the square tiling $\{4,4\}$, the triangular tiling $\{3,6\}$, and the hexagonal tiling $\{6,3\}$. In contrast, tilings of the hyperbolic plane corresponds to the inequality $1/p+1/q<1/2$, which allows for an infinite number of regular tessellations.
In the main body of this work,we shall focus on the class of tessellations with $q>3$. As Sec. V.2 will discuss, models on hyperbolic tessellations with $q=3$ are strongly constrained and have a different behavior from other tesselation. The construction of our lattices follows a set of inflation rules [32], growing the lattice layer-by-layer ($k=1,2,\ldots$) from a central polygon, denoted $\sigma$, as depicted in Figure 1. To formalize the construction, we classify the polygons and vertices at each layer based on their connectivity to the preceding layer:
@@
$\beta$ $\rightarrow\alpha^{r}(\beta\alpha^{q-3})^{p-4}\beta\alpha^{r}$
@@ IV.1 Plaquette Ising Model
To establish the framework for the hyperbolic fracton model, we first review the foundational concepts of fracton physics through the classical toy model: the Plaquette Ising Model (PIM).
@@ IV.2 Hyperbolic Fracton Model
The Hyperbolic Fracton Model, originally introduced for the $\{5,4\}$ tessellation [16, 17], generalizes the plaquette Ising model to generic $\{p,q\}$ hyperbolic tessellations. As in the PIM, Ising spins $S_{i}^{z}$ reside at the center of each $p$-gon and $q$ polygons meet at vertices. The Hamiltonian (18) retains its form but vertex operators extend to any $q>3$ spins (Fig. 4).
In the original HFM model [16, 17], geodesics and their associated subsystem symmetries play a fundamental role in establishing holographic properties. On the Euclidean lattice, geodesics correspond to straight lines along the square lattice edges. Meanwhile, in the hyperbolic $\{5,4\}$ lattice geodesics manifest as circular arcs that orthogonally intersect the boundary (Fig. 4a). These subsystem symmetries were used to demonstrate that the HFM exhibits several key properties of holography. First, the model realizes the Rindler reconstruction, which states that spin configurations on any connected boundary subregion determine bulk states within their minimal geodesic wedge (the discrete analog of an entanglement wedge). Second, the mutual information for bipartitions of connected boundary subregions satisfies the Ryu-Takayanagi formula. Finally, a naively defined black hole in the model has an entropy that follows the Bekenstein-Hawking formula.
In hyperbolic lattices with $q>4$, the subsystem symmetries become much more complicated. For odd $q$, there are no geodesics formed by the edges of the lattice, and for even $q$, the geodesics do not define subsystem symmetry directly [35]. The lack of systematic construction of subsystem symmetries motivate our generalization of the HFM through the use of inflation rules (Section III), as they provide an alternative to construct subsystem symmetries and demonstrate holographic properties. Crucially, in section V we demonstrate that the original HFM exhibits an extensive entropy that scales with the system’s volume. This stands in contrast with Euclidean fracton models, where subsystem symmetries yield a sub-extensive degeneracy.
@@ V.1 General Tessellations
Let us now compute the ground state degeneracy (GSD) of a general HFM. We shall respectively define $N_{\alpha}^{k}$, $N_{\beta}^{k}$ and $N_{p}^{k}$, the number of $\alpha$ and $\beta$ polygons and the total number of polygons (i.e., spins) on layer $k$. $N_{v}^{k}$ is the number of vertices between layers $k$ and k1k-1. By definition for $k\geq 1$,
@@
Another way to derive the above equation is the vertex representation where we subtract the number of constraints per layers. Since the number of degrees of freedom (DOF) in a layer is the total number of spins, $N_{p}^{k}$, and the number of constraints imposed by each vertex is $N_{v}^{k}$, the number of free DOF per layer is $N_{p}^{k}-N_{v}^{k}=N^{k}_{\alpha}$ for $k\geq 1$ (see Eqs. (19)).
[h]
The residual ground-state entropy is
@@
Once normalized by the total number of spins $N_{p}=\sum_{k=1}^{l}(N^{k}_{\alpha}+N^{k}_{\beta})+1$, we have in the thermodynamic limit:
@@
A non-zero residual entropy indicates that the system has an extensive degeneracy, as a consequence of the exponential growth of the hyperbolic lattice. This result stands in contrast to Euclidean fracton models whose sub-extensive degeneracy yields a residual entropy of zero.
@@ V.2 The $\{p,3\}$ Tessellations
The case where $q=3$ represents a highly constrained limit of the HFM. In this geometry, each vertex is shared by exactly three polygons, which fundamentally alters the recursive structure of the lattice and the nature of its ground state degeneracy. In particular, the inflation rules are distinct due to the different local connectivity. The polygon types are now denoted $\sigma$ (central), $\beta$ (sharing an edge with the previous layer), and $\gamma$ (sharing a vertex). The inflation rules, $\tau$, are given by:
@@
$\gamma$ $\to\gamma^{1/2}\beta^{p-6}\gamma^{1/2}$
@@
  • p is even: For even values of $p$, the ground state is four-fold degenerate. This degeneracy arises from two distinct binary choices: One degree of freedom corresponds to the orientational choice of the central spin and a second degree of freedom arises from the choice of the initial spin orientation on the first layer ($k=1$). This is illustrated in Figure 6(ad)(a-d).
V.3 Subsystem Symmetries

### v2

@@ Abstract
We generalize the Hyperbolic Fracton Model from the $\{5,4\}$ tessellation to generic {p,qp,q} tessellations, and investigate its core properties – ground-state degeneracy, fracton mobility, and holographic correspondence – bringing to light a richer and more intricate structure than the original case. The ground-state degeneracy and subsystem symmetries are computed exactly through the inflation rule, but do not admit a simple, symmetric pattern. For $q=3$, we find that the degeneracy is finite, either unique or four-fold. For the {4,4} square lattice as the flat limit of $q=4$, the degeneracy is well known to be subextensive. But for all other tessellations of $q>3$, the degeneracy becomes extensive, including for the {3,6} lattice on flat space which maps onto the honeycomb color code. The fracton excitation number follows exponential-in-distance and algebraic-in-lattice-size growing patterns when moving outward, and depends sensitively to the tessellation geometry, differing qualitatively from both type-I or type-II fracton model on flat lattices. Despite this increased complexity, the hallmark holographic features — subregion duality via Rindler reconstruction, the Ryu-Takayanagi formula for mutual information, and effective black hole entropy scaling with horizon area — remain valid. These results demonstrate that the holographic correspondence in fracton models persists in generic tessellations, and provide a natural platform to explore more intricate subsystem symmetries and fracton physics.
@@ I Introduction
Fracton phases of matter have emerged as an exciting extension of conventional paradigms in condensed matter physics and quantum information theory. Fracton systems host quasiparticle excitations with fundamentally restricted mobility, ranging from immobile fractons to subdimensional particles confined to lines or planes [1, 2, 3, 4, 5, 6, 7], and motivate a wide range of research in quantum error correction codes, new phases of matter, and constrained dynamics. These unconventional properties reflect an intimate interplay between many-body entanglement, lattice geometry, and generalized symmetries acting on lower-dimensional subsystems [8, 9, 10].
A distinctive feature of fracton physics is its sensitivity to lattice geometry beyond topology alone. Unlike conventional topological phases whose universal properties are largely insensitive to lattice details, fracton models can exhibit qualitatively different behavior when formulated on lattices with different coordination, curvature, or dimensional embedding. This geometric dependence has led to fruitful connections with elasticity theory, higher-rank gauge theories, and, in certain limits, structures reminiscent of gravity [11, 12, 13]. As a result, fracton systems provide a natural arena in which concepts from quantum information and condensed matter intersect with ideas traditionally associated with high-energy physics.
In parallel, the holographic principle – most prominently realized in the anti-de Sitter/conformal field theory (AdS/CFT) correspondence [14, 15, 16] – has inspired a broad search for simplified, non-string-theoretic toy models that capture selected aspects of holography [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. While fracton models are not gravitational theories, their constrained dynamics, subregion symmetries, and sensitivity to geometry make them appealing candidates for exploring holography-inspired phenomena in a controlled many-body setting [29, 30, 31].
Motivated by these perspectives, in Refs. [29, 30, 31] we introduced and studied the Hyperbolic Fracton Model (HFM), the simplest classical fracton model defined on a regular hyperbolic lattice, specifically the $\{5,4\}$ tessellation. Hyperbolic lattices provide a discrete realization of negatively curved geometry and naturally implement a hierarchical bulk-boundary structure [32]. Despite its simplicity, the HFM exhibits a number of remarkable properties, including extensive ground-state degeneracy, emergent boundary degrees of freedom, and geodesic-like constraints that encode bulk-boundary relations. Subsequent work further clarified these features and established connections to $p$-adic and tensor-network-inspired toy models of holography [33, 34].
@@
Here we generalize the original classical HFM to all $\{p,q\}$ tessellations of the hyperbolic plane, extending beyond the original simple case, and discover unconventional behavior without analogue in the type-I and type-II fracton models in flat 3D lattices [35]. The construction of our model is based on a set of geometry-dependent, recursive inflation rules that dictate the subsystem symmetries of the system. Through a detailed analysis of the model’s ground state degeneracy and properties, we derive scaling laws governed by the tessellations’ parameters $p$ and $q$. We find that almost all tessellations with $q>3$ support an extensive ground-state degeneracy. In particular the HFM on the Euclidean {3,6} lattice is equivalent to the classical limit of the honeycomb color code with a $\mathbb{Z}_{2}$ gapped topological order phase [36]. Another central result of our generalized model is the Rindler reconstruction, a procedure through which bulk spin configurations within a minimal wedge are uniquely determined by measurements on a corresponding boundary subregion. This provides a discrete analogue to the entanglement wedge reconstruction in continuum AdS/CFT. We also demonstrate that the mutual information between complementary boundary subregions, which is the classical analogue of entanglement entropy, follows a discrete form of the Ryu-Takayanagi formula [37], therefore linking boundary entanglement to the length of the shared boundary. Furthermore, we establish that the removal of bulk spins simulates the formation of a black hole. This induces a change in entropy proportional to the boundary area of the defect region, and offers a discrete analogue of the Bekenstein-Hawking entropy formula [38]. Finally, we have made quantitative analysis of how fractons grow exponentially in distance and algebraically in system size when pushed away to outwards layers. This feature is qualitatively different from both type-I and type-II fracton models on flat cubic lattice, and applies to quantum fracton orders defined on lattices embedded in 2D hyperbolic space $\times$ 1D flat space too [39].
@@
This paper is structured as follows: Section II provides a brief review of the essential concepts of the AdS/CFT correspondence that underpin our model. Section III elaborates on the recursive construction of generic hyperbolic tessellation geometry. Section IV introduces the Plaquette Ising model as a foundation for the hyperbolic generalization and presents the Hyperbolic Fracton Model [29, 30]. The following sections present our results. In Section V, we derive exact expressions for the ground state degeneracy and residual entropy. Then, the holographic properties of the model are elaborated through the Rindler reconstruction in Section VI, mutual information calculations in Section VII, and an analysis of black hole entropy in Section VIII. Stepping away from the ground-state properties, section IX is devoted to the analysis of fracton excitations. Finally we conclude our paper with a discussion in Section X.
@@ II AdS/CFT correspondence
The first concrete realization of the holographic principle is the AdS/CFT correspondence [14], which is a conjectured duality between a $(d+1)$-dimensional gravitational theory in an Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) on its boundary. AdS space is a maximally symmetric Lorentzian manifold with negative curvature. Meanwhile, a CFT is a field theory that is invariant under conformal symmetries, particularly powerful for describing systems at their critical points. The AdS/CFT duality is a strong-weak duality, where the boundary CFT that is strongly coupled and analytically intractable, corresponds to a weakly coupled gravitational theory in the bulk.
@@
In our work, we will focus on a constant time slice of AdS3, which has the geometry of a hyperbolic plane $\mathbb{H}^{2}$ with constant, negative curvature.
@@ III Hyperbolic Tessellations
A regular tessellation is a tiling of a surface by identical, regular polygonal faces, where the same number of faces meet at each vertex. These structures are uniquely classified by their Schläfli symbol $\{p,q\}$, where $p$ is the number of edges of each polygonal face, and $q$ is the number of faces meeting at each vertex of the tiling. The geometry of the embedding space is determined by the relationship between $p$ and $q$. For a flat Euclidean plane, tessellations must satisfy $1/p+1/q=1/2$, which permits only three configurations: the square tiling $\{4,4\}$, the triangular tiling $\{3,6\}$, and the hexagonal tiling $\{6,3\}$. In contrast, tilings of the hyperbolic plane negatively curved space corresponds to the inequality $1/p+1/q<1/2$, which allows for an infinite number of regular tessellations.
All $\{p,q\}$ tessellations shown in this work are drawn using the conformal Poincaré disk model of the hyperbolic plane $\mathbb{H}^{2}$, consistent with the representations used in earlier papers. In this model, hyperbolic geodesics are represented by circular arcs orthogonal to the boundary of the disk. The apparent straight edges in some figures do not correspond to a change of hyperbolic representation (e.g., to the Klein model) – it is merely for better visualization.
In the main body of this work, we shall focus on the class of tessellations with $q>3$. As we will discuss in section V.2, models with $q=3$ are strongly constrained by the trigonal symmetry of the vertices and have a different behavior from other tessellation.
The construction of our lattices for $q>3$ follows a set of inflation rules [32], growing the lattice layer-by-layer ($k=1,2,\ldots$) from a central polygon, denoted $\sigma$, as depicted in Figure 1. To formalize the construction, we classify the polygons and vertices at each layer based on their connectivity to the preceding layer:
@@
$\beta$ $\rightarrow\alpha^{r}(\beta\alpha^{q-3})^{p-4}\beta\alpha^{r}$ (1)
@@ IV.1 Plaquette Ising Model
To establish the framework for the hyperbolic fracton model, we first review the foundational concepts of fracton physics through the classical toy model: the Plaquette Ising Model (PIM) on the {4,4} tessellation.
@@ IV.2 Hyperbolic Fracton Model
The Hyperbolic Fracton Model, originally introduced for the $\{5,4\}$ tessellation [29, 30], generalizes the plaquette Ising model to generic $\{p,q\}$ hyperbolic tessellations. As in the PIM, classical Ising spins $S_{i}^{z}$ reside at the center of each $p$-gon and $q$ polygons meet at vertices. The Hamiltonian (19) retains its form but vertex operators extend to $q$ spins (Fig. 4).
In the original HFM model [29, 30], geodesics and their associated subsystem symmetries play a fundamental role in establishing holographic properties. On the Euclidean lattice, geodesics correspond to straight lines along the square lattice edges. Meanwhile, in the hyperbolic $\{5,4\}$ lattice geodesics manifest as circular arcs that orthogonally intersect the boundary (Fig. 4a). These subsystem symmetries were used to demonstrate that the HFM exhibits several key properties of holography [29, 34]. First, the model realizes the Rindler reconstruction, which states that spin configurations on any connected boundary subregion determine bulk states within their minimal geodesic wedge (the discrete analog of an entanglement wedge). Second, the mutual information for bipartitions of connected boundary subregions satisfies the Ryu-Takayanagi formula. Finally, a naively defined black hole in the model has an entropy that follows the Bekenstein-Hawking formula.
In hyperbolic lattices with $q>4$, the subsystem symmetries become more complex. For odd $q$, there are no geodesics formed by the edges of the lattice, and for even $q$, the geodesics do not define subsystem symmetry directly [33]. The lack of systematic construction of subsystem symmetries motivate our generalization of the HFM through the use of inflation rules (Section III), as they provide an alternative to construct subsystem symmetries and demonstrate holographic properties. Crucially, in section V we demonstrate that the original HFM exhibits an extensive entropy that scales with the system’s volume. This stands in contrast with the Plaquette Ising Model, where subsystem symmetries yield a sub-extensive degeneracy.
@@ V.1 General Tessellations
Let us now compute the classical ground state degeneracy (GSD) of a general HFM. We shall respectively define $N_{\alpha}^{k}$, $N_{\beta}^{k}$ and $N_{p}^{k}$, the number of $\alpha$ and $\beta$ polygons and the total number of polygons (i.e., spins) on layer $k$. $N_{v}^{k}$ is the number of vertices between layers $k$ and k1k-1. By definition for $k\geq 1$,
@@
Another way to derive the above equation is the vertex representation where we subtract the number of constraints per layers. Since the number of degrees of freedom (DOF) in a layer is the total number of spins, $N_{p}^{k}$, and the number of constraints imposed by each vertex is $N_{v}^{k}$, the number of free DOF per layer is $N_{p}^{k}-N_{v}^{k}=N^{k}_{\alpha}$ for $k\geq 1$ (see Eqs. (20)). The ground-state entropy is
@@
Once normalized by the total number of spins $N_{p}=\sum_{k=1}^{l}(N^{k}_{\alpha}+N^{k}_{\beta})+1$ and taking the thermodynamic limit, we get the residual ground-state entropy:
@@
This result is valid for all tessellations {p,q>3p,q>3}. $R_{\infty}$ is always finite, except for $p=q=4$ where it diverges. This exception corresponds to the well-known sub-extensive entropy of the Plaquette Ising Model on the Euclidean square lattice. The HFM on negatively curved space with $q>3$ thus always supports an extensive degeneracy, in sharp contrast with the PIM.
@@ V.2 The $\{p,3\}$ Tessellations
The case where $q=3$ represents a constrained limit of the HFM. In this geometry, each vertex is shared by exactly three polygons, which alters the recursive structure of the lattice and the nature of its ground state degeneracy. In particular, the inflation rules are distinct due to the different local connectivity. The polygon types are now denoted $\sigma$ (central), $\beta$ (sharing an edge with the previous layer), and $\gamma$ (sharing three vertices with the previous layer). The inflation rules, $\tau$, are given by:
@@
$\gamma$ $\to\gamma^{1/2}\beta^{p-6}\gamma^{1/2}$ (25)
@@
  • p is even: The ground state is four-fold degenerate. This degeneracy arises from two distinct binary choices: One degree of freedom corresponds to the orientational choice of the central spin and a second degree of freedom arises from the choice of the initial spin orientation on the first layer ($k=1$). This is illustrated in Figure 6(ad)(a-d).
V.3 Tessellations in flat space
V.3.1 Square lattice ({4,4} tessellation)
Although the Plaquette Ising model on the square lattice is on a flat space, its inflation rule is given by the same formula (1) as for hyperbolic lattices. It means one can directly apply Eq. (22) to obtain the ground-state entropy,
$S=k_{B}\log{2}\times\left(1+\sum_{k=1}^{l}4\right)=k_{B}\log{2}\times\left(1+4l\right).$
Since a {4,4} tessellation of $l$ layers has $(2l+1)^{2}$ spins, the intensive entropy vanishes in the thermodynamic limit as expected,
$s=\lim_{l\to\infty}k_{B}\log{2}\times\cfrac{1+4l}{(1+2l)^{2}}\to 0.$
On the other hand, since we have $N_{\text{surf}}=8l$ spins on the boundary, the ground-state entropy scales linearly with the boundary area
$\lim_{l\to\infty}\cfrac{S}{N_{\text{surf}}}=\lim_{l\to\infty}k_{B}\log{2}\times\left(\cfrac{1+4l}{8l}\right)=\frac{1}{2}k_{B}\log{2}.$
as expected from the discussion of section IV.1.
V.3.2 Hexagonal lattice ({6,3} tessellation)
The inflation rule (25) also applies to the $\{6,3\}$ tessellation on flat space
$\tau:\quad\sigma$ $\to\beta^{6}$
$\beta$ $\to\gamma^{1/2}\beta\gamma^{1/2}$
$\gamma$ $\to\gamma.$
Hence, as for the other $\{p,3\}$ tessellations with even values of $p$, the ground state degeneracy is four fold, as illustrated in Fig. 7. Since $\gamma$ spins are in contact with two spins in the previous layer, their orientations are fixed by the previous layer ($q=3$). And since $\beta$ spins are surrounded by $\gamma$ spins in layer $l$, their orientations are also fixed. It means the finite degeneracy of $\{p,3\}$ tessellations is attached to the local trigonal symmetry of vertices, and not to the negative space curvature.
Figure 7: The four ground states of the hexagonal tessellation {6,3}, following the same reasoning as in Fig. 6 for {p,3p,3} hyperbolic tessellations with even values of $p$.
V.3.3 Triangular lattice ({3,6} tessellation)
Since the inflation formula of Eq. (1) is valid for the $\{3,6\}$ tessellation, one can directly use the formula of the ground-state degeneracy (23), which gives
$s=\lim_{l\to\infty}\cfrac{S}{N_{p}}\to\frac{1}{2}k_{B}\log{2},$
with $R_{\infty}=1$. The extensive entropy can be understood locally. With six triangles meeting at a vertex the system is under-constrained and allows for loop fluctuations in the ground state, as illustrated in Fig. 8. Similar loop fluctuations are allowed for all hyperbolic $\{3,q\}$ tessellations.
Figure 8: Example of loops of spin flips that can be formed on the $\{3,6\}$ tessellation, contributing to the extensive ground-state degeneracy. All loops can be built from the elementary hexagon loops.
This model is equivalent to the classical limit of the honeycomb color code, which realizes a $\mathbb{Z}_{2}$ gapped topological order phase. The local symmetry (flipping a loop) corresponds to exactly the dynamics of the honeycomb color code [36]. The {3,6} tessellation thus offers the only extensive degeneracy of the HFM in Euclidean space.
As a running summary, the degeneracy of the HFM is essentially linked to the connectivity of the lattice; finite for $q=3$, extensive for $q>3$, with the {4,4} Plaquette Ising Model as a marginal case with subextensive degeneracy.
V.4 Subsystem Symmetries